OpenStudy (anonymous):

Jade hiked 1 1/4 mile on the trail. Gretta hiked 1 1/2 times as far on the trail as Jade. How far did Gretta hike on the trail?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@litzyceron

OpenStudy (anonymous):

\[\large \sf 1\frac{1}{4}=\frac{5}{4}\]\[\large \sf 1\frac{1}{2}=\frac{3}{2}\] \[\large \sf \frac{5}{4} \times \frac{3}{2} =?\]

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

Yeah, although you could put that into a mixed fraction form to make it look better

OpenStudy (anonymous):

so it would be 1 7/8

OpenStudy (anonymous):



OpenStudy (bubblegum.):

\(\Huge\color{green} ✓ \)

OpenStudy (anonymous):

thank u

OpenStudy (bubblegum.):

no problem :)